# Multi-Scale Fusion for Object Representation

### v2

@@ 1 Introduction
According to vision cognition research, humans perceive a visual scene as objects and relations among them for higher-level cognition such as understanding, reasoning, prediction, planning and decision-making (Bar, 2004; Cavanagh, 2011; Palmeri & Gauthier, 2004). In computer vision, Object-Centric Learning (OCL) is a promising way to achieving similar effects – Images or video frames are represented as sparse object-level feature vectors, known as slots, rather than dense (super-)pixel-level feature maps under weak or self-supervision (Greff et al., 2019; Burgess et al., 2019). Meanwhile, segmentation masks of objects and the background are generated as byproducts, intuitively reflecting the corresponding object representation quality of these slots.
Object-level representation learning is highly influenced by textures. Early OCL methods, such as the mixture-based (Locatello et al., 2020; Kipf et al., 2022; Elsayed et al., 2022), which directly reconstruct the input pixels as supervision, often struggle with objects that have complex textures. To address this issue, mainstream OCL methods, such as transformer-based (Singh et al., 2022a; c), foundation-based (Seitzer et al., 2023; Zadaianchuk et al., 2024) and diffusion-based (Jiang et al., 2023; Wu et al., 2023b), leverage well-learnt intermediate representation for guided reconstruction, so as to drive slots to aggregate as much object-level information as possible.
@@
Objects appear at different scales in images or videos due to imaging distances and actual sizes. Encoders/decoders usually perform well on certain pattern scales, necessitating ”divide-and-rule” (Zou et al., 2023; Minaee et al., 2021). Mainstream OCL, however, overlooks such an issue.
@@ 2 Related Work
VAE & OCL. Pretrained VAE (Im Im et al., 2017; Van Den Oord et al., 2017), like dVAE or VQ-VAE, can provide representations that highlight object separability in the input by suppressing texture redundancies for OCL. With such guidance, transformer-based OCL methods (Singh et al., 2022a; c), like SLATE and STEVE, decode slots, which are extracted by SlotAttention (Locatello et al., 2020) from input images or video frames, into the input via a transformer decoder. Similarly, diffusion-based methods (Jiang et al., 2023; Wu et al., 2023b), like SlotDiffusion, decode slots via a diffusion model. Slots are driven to aggregate as much object information for good object representation. We improve these two lines of work. Foundation-based methods (Seitzer et al., 2023; Zadaianchuk et al., 2024), like DINOSAUR, are also included for complete comparison.
Channel Grouping. Widely adopted in CNNs (Zhang et al., 2018) and ViTs (Gu et al., 2022), grouping features along the channel dimension helps learn expressive representations. Grouping can be performed on features (Krizhevsky et al., 2012; Huang et al., 2018; Chen et al., 2019), or on weights (Zhao et al., 2021; 2022). SysBinder (Singh et al., 2022b) explores this in OCL by grouping slots queries to aggregate different object information, for emerging attributes under extra supervision. The most recent GDR (Zhao et al., 2024) groups the VAE codebook from features into combinatorial attributes, achieving better generalization and convergence in object representing. Our work also groups it but for scale-invariance and -variance.
Multi-Scale. Objects exist at different scales due actual sizes and imaging distances. CNN (Zhang et al., 2018) and ViT (Gu et al., 2022) typically have their comfort zones as the scale distribution of objects in training data is always nonuniform. In computer vision, like detection (Zou et al., 2023) and segmentation (Minaee et al., 2021), the multi-scale issue is significant, leading to the wide adoption of “divide-and-rule”. The image pyramid (Adelson et al., 1984) produces multi-scale representations by resizing; Feature pyramid network (Lin et al., 2017) and variants mix features of different layers by channel-concat and element-wise-sum. We are the first to realize multi-scale fusion on VAE, augmenting every scale with the high-quality representation in other scales.
@@ 3 Proposed Method
We propose Multi-Scale Fusion (MSF), to explicitly address the multi-scale issue in VAE intermediate representation, and thereby to guide Object-Centric Learning (OCL) better. Our technique is applicable to both transformer-based OCL (Singh et al., 2022a; c) classics and diffusion-based OCL (Wu et al., 2023b) state-of-the-arts, which are our basis methods. Note that two types of methods use dVAE (Im Im et al., 2017) and VQ-VAE (Van Den Oord et al., 2017) respectively, but we unify them with VQ-VAE like that in GDR (Zhao et al., 2024).
@@ 3.1 Background: OCL with VAE Guidance
Initially, the primary encoder encodes the input into a feature map. $\bm{\phi}_{\mathrm{e}}:\bm{T}\rightarrow\bm{F}$, where $\bm{F}\in\mathbb{R}^{h\times w\times c}$, and $\bm{\phi}_{e}$ is a CNN (He et al., 2016) or ViT (Caron et al., 2021) module. Typically, $h=w=32$ and $c=256$.
Then, the Slot Attention (SA) (Locatello et al., 2020) is used to aggregate the dense feature map $\bm{F}$ into sparse object-level feature vectors, i.e., slots $\bm{S}\in\mathbb{R}^{s\times c}$, along with the byproduct, i.e., segmentation masks $\bm{M}\in\mathbb{R}^{s\times h\times w}$, under query vectors $\bm{S}_{0}$. $\bm{\phi}_{\mathrm{sa}}:(\bm{S}_{0},\bm{F})\rightarrow(\bm{S},\bm{M})$, which is basically an iterative Query-Key-Value attention (Bahdanau et al., 2015) mechanism with $\bm{S}_{0}$ as the query and $\bm{F}$ as the key and value. Typically, $s$ is a predefined number, usually the maximum number of objects plus the background in an image or video sample of a dataset.
Meanwhile, pretrained VAE (Im Im et al., 2017; Van Den Oord et al., 2017) converts $\bm{T}$ into continuous intermediate representation. $\bm{\phi}_{\mathrm{e}}^{\mathrm{v}}:\bm{T}\rightarrow\bm{Z}$, where $\bm{Z}\in\mathbb{R}^{h\times w\times c}$ and typically $h=w=32$ and $c=256$. This continuous representation is quantized into discrete representation by matching continuous representation with codebook and selecting matched codes to form discrete representation. $f_{\mathrm{q}}=f_{\mathrm{m}}\circ f_{\mathrm{s}}:(\bm{Z},\bm{C})\rightarrow(\bm{P},\bm{I},\bm{X})$, where codebook $\bm{C}\in\mathbb{R}^{m\times c}$, matching probabilities $\bm{P}\in\mathbb{R}^{h\times w\times m}$, matched indexes $\bm{I}\in\mathbb{R}^{h\times w}$ and discrete representation $\bm{X}\in\mathbb{R}^{h\times w\times c}$. The matching $f_{\mathrm{m}}$ and selection $f_{\mathrm{s}}$ are implemented as below, respectively
$\bm{P}=\mathrm{softmax}_{m}(-||\bm{Z}-\bm{C}||_{2}),\quad\bm{I}=\mathrm{argmax}_{m}(\bm{P})$ (1)
@@
Lastly, with $\bm{S}$ as the condition, the primary decoder reconstructs the input guided by VAE discrete intermediate representation conditioned on $\bm{S}$, which drives every slot to aggregate as much object information as possible. $\bm{\phi}_{\mathrm{d}}:(\bm{X},\bm{S})\rightarrow\bm{X}^{\prime}$, where $\bm{X}^{\prime}\in\mathbb{R}^{h\times w\times c}$. For transformer-based OCL, a transformer decoder (Vaswani et al., 2017) with internal masking performs the reconstruction as classification. For diffusion-based OCL, a diffusion model, typically a conditional UNet (Rombach et al., 2022), carries out the reconstruction as regression of the noise added to $\bm{X}$.
@@ 3.2 Naive Multi-Scale Training
There can be objects of different sizes in images or videos, while models usually have best-at patterns scales due to non-uniform data distributions. This makes image pyramid (Adelson et al., 1984) an effective technique, where the input is resized into different sizes and processed in parallel. Such multi-scale representations make the pattern scale distribution more uniform for training, and ensures different sized objects to fall within models’ comfort zone for testing.
Input image or video frame $\bm{T}_{1}:=\bm{T}$ is down-sampled with typical bi-linear interpolation into $\bm{T}_{n}\in\mathbb{R}^{\frac{h_{0}}{2^{n-1}}\times\frac{w_{0}}{2^{n-1}}\times c_{0}}$, where $n=1...N$ means different spatial scales and we use $N=3$.
@@ 3.3 Inter/Intra-Scale Quantized Fusion
Firstly, we project VAE continuous representation of a certain scale $\bm{Z}_{n}$ into two copies, $\bm{Z}_{n}^{\mathrm{si}}$ for scale-invariance and $\bm{Z}_{n}^{\mathrm{sv}}$ for scale-variance, respectively. These two will finally be combined back, thus we employ the invertible project-up-project-down design proposed in Zhao et al. (2024) to maintain the inductive biases added in between.
@@
Meanwhile, for the scale-invariant continuous representation, we conduct inter-scale fusion over scale-invariant quantization of all scales. $f_{\mathrm{rf}}:\{(\bm{Z}_{n}^{\mathrm{si}},\bm{C}^{\mathrm{si}})\}\rightarrow\{(\dot{\bm{P}}_{n}^{\mathrm{si}},\dot{\bm{I}}_{n}^{\mathrm{si}},\dot{\bm{X}}_{n}^{\mathrm{si}})\}$, where $n=1...N$. The shared codebook enforces scale-invariant representation learning.
Now that all scales $\{\bm{X}_{n}^{\mathrm{si}}\}$ consist of codes in a shared codebook $\bm{C}^{\mathrm{si}}$, matching probabilities $\{\bm{P}_{n}^{\mathrm{si}}\}$ for the same (super-)pixel region under different scales indicate how confident the model is about the representation quality. We utilize this to augment the low-quality representation in some scales with the high-quality in other scales. Take scales $n-1$, $n$ and $n+1$ as an example:
@@
    Given any super-pixel $\bm{z}_{n}^{\mathrm{si}}:=\bm{Z}_{n}^{\mathrm{si},(i,j,:)}\in\mathbb{R}^{c}$ in any scale $n$, find the corresponding region/super-pixel(s) in other scales, and mean them along all spatial dimensions as voting. For scale $n-1$, we choose $\bm{Z}_{n-1}^{\mathrm{si},(2i:2i+2,2j:2j+2,:)}\in\mathbb{R}^{2\times 2\times c}$ and mean it into a vector $\bm{z}_{n-1}^{\mathrm{si}}\in\mathbb{R}^{c}$; For scale $n+1$, we choose $\bm{z}_{n+1}^{\mathrm{si}}:=\bm{Z}_{n+1}^{\mathrm{si},(\lfloor\frac{i}{2}\rfloor,\lfloor\frac{j}{2}\rfloor,:)}\in\mathbb{R}^{c}$.
@@ 3.4 Resource Consumption Analysis
Codebook-related parameters. The basis methods use a codebook of size $(m,c)=4096\times 256$, which amounts to $1.048$M. We use (1) invertiable projection to project the continuous representation into two copies (and project them back), up to $c\times 2c=2^{17}$; (2) four groups of codebooks of size $(m^{0.5},c)=(64,256)$, one shared for $N=3$ scale-invariant scales, three specifically for $N=3$ scale-variant scales, up to $m^{0.5}\times c\times(1+N)=2^{16}$. Thus our codebook-related parameters are $0.196$M – we have $81.25\%$ fewer codebook parameters than the basis methods.
Computation and memory consumption. The exact calculation relates with the VAE encoder/decoder layers, like Conv2d, GroupNorm, Mish, etc., thus has too many details. But due to the $O(n^{2})$ complexity nature of convolution layers, we can quickly get the estimation in ratio. Denote the computation of the original VAE as unit, then our $N=3$ scales’ computations would be $1$, $\frac{1}{4}$ and $\frac{1}{16}$, since the latter two has 2×2$\times$2 × and 4×4$\times$4 × down sampling rates. Thus in VAE pretraining, our technique requires roughly $31.25\%$ more computation. The memory consumption increase is similar.
@@ 4.1 OCL Performance
To evaluate the quality of OCL object representations, i.e., slots, we use the accuracy of OCL’s byproduct (unsupervised) segmentation as an intuitive measurement. Metrics including ARI (Adjusted Rand Index) [^1] https://scikit-learn.org/stable/modules/generated/sklearn.metrics.adjusted_rand_score.html, ARIfg (foreground), mIoU (mean Intersection-over-Union) [^2] https://scikit-learn.org/stable/modules/generated/sklearn.metrics.jaccard_score.html and mBO (mean Best Overlap) (Caron et al., 2021) are used to measure OCL’s byproduct segmentation accuracy. Datasets are either synthetic images, i.e., ClevrTex [^3] https://www.robots.ox.ac.uk/~vgg/data/clevrtex/; or real-world images, i.e., COCO [^4] https://cocodataset.org/#panoptic-2020 and VOC [^5] http://host.robots.ox.ac.uk/pascal/VOC/voc2012/index.html; or synthetic videos, i.e., MOVi-C/D/E [^6] https://github.com/google-research/kubric/tree/main/challenges/movi.
To evaluate how MSF boosts OCL, we use transformer-based classics, SLATE (Singh et al., 2022a) for images and STEVE (Singh et al., 2022c) for videos; use diffusion-based state-of-the-arts, SlotDiffusion (Wu et al., 2023b) for images. The foundation-based DINOSAUR (Seitzer et al., 2023) and competitors SysBinder (Singh et al., 2022b) and GDR (Zhao et al., 2024) are compared too. All primary encoders are unified as DINO (Caron et al., 2021) to form strong baselines.
Our MSF is a general improver to both transformer-based classics and diffusion-based state-of-the-arts, as shown in Tab. 2, 2 and 4. On complex synthetic dataset ClevrTex, MSF improves SLATE significantly, up to 12% in ARIfg. On challenging real-world dataset COCO, MSF still manages to improve SLATE by 6% in ARIfg. MSF even boosts SlotDiffusion, which is state-of-the-art, by 3% in ARI on VOC. Compared with other improvers, i.e., SysBinder@$g$2 and GDR@$g$2, our MSF also beats or at least ties them on SLATE, while beats them consistently on STEVE. On SlotDiffusion, our MSF beats the strong competitor GDR@$g$2, and defeats DINOSUAR.
By the way, our MSF boosts OCL performance on COCO and VOC less than on other datasets. We attribute this to those object sizes are often out of the coverage of scales 1~3. Namely, many humans, trees and grass lands in those real-world images take up the whole scene, and either down sampling 4$\times$ (scale 1) or 16$\times$ (scale 3) cannot produce scale-invariant patterns for inter-scale fusion.

### v3

@@ 1 Introduction
According to vision cognition research, humans perceive a visual scene as objects and relations among them for higher-level cognition such as understanding, reasoning, prediction, planning and decision-making (bar2004visual; cavanagh2011visual; palmeri2004visual). In computer vision, Object-Centric Learning (OCL) is a promising way to achieving similar effects – Images or video frames are represented as sparse object-level feature vectors, known as slots, rather than dense (super-)pixel-level feature maps under weak or self-supervision (greff2019iodine; burgess2019monet). Meanwhile, segmentation masks of objects and the background are generated as byproducts, intuitively reflecting the corresponding object representation quality of these slots.
Object-level representation learning is highly influenced by textures. Early OCL methods, such as the mixture-based (locatello2020slotattent; kipf2021savi; elsayed2022savipp), which directly reconstruct the input pixels as supervision, often struggle with objects that have complex textures. To address this issue, mainstream OCL methods, such as transformer-based (singh2021slate; singh2022steve), foundation-based (seitzer2023dinosaur; zadaianchuk2024videosaur) and diffusion-based (jiang2023lsd; wu2023slotdiffuz), leverage well-learnt intermediate representation for guided reconstruction, so as to drive slots to aggregate as much object-level information as possible.
@@
Objects appear at different scales in images or videos due to imaging distances and actual sizes. Encoders/decoders usually perform well on certain pattern scales, necessitating ”divide-and-rule” (zou2023objectdetection; minaee2021imagesegmentation). Mainstream OCL, however, overlooks such an issue.
@@ 2 Related Work
VAE & OCL. Pretrained VAE (im2017dvae; van2017vqvae), like dVAE or VQ-VAE, can provide representations that highlight object separability in the input by suppressing texture redundancies for OCL. With such guidance, transformer-based OCL methods (singh2021slate; singh2022steve), like SLATE and STEVE, decode slots, which are extracted by SlotAttention (locatello2020slotattent) from input images or video frames, into the input via a transformer decoder. Similarly, diffusion-based methods (jiang2023lsd; wu2023slotdiffuz), like SlotDiffusion, decode slots via a diffusion model. Slots are driven to aggregate as much object information for good object representation. We improve these two lines of work. Foundation-based methods (seitzer2023dinosaur; zadaianchuk2024videosaur), like DINOSAUR, are also included for complete comparison.
Channel Grouping. Widely adopted in CNNs (zhang2018shufflenet) and ViTs (gu2022hrvit), grouping features along the channel dimension helps learn expressive representations. Grouping can be performed on features (krizhevsky2012alexnet; huang2018condensenet; chen2019octconv), or on weights (zhao2021lhc; zhao2022coc). SysBinder (singh2022sysbind) explores this in OCL by grouping slots queries to aggregate different object information, for emerging attributes under extra supervision. The most recent GDR (zhao2024gdr) groups the VAE codebook from features into combinatorial attributes, achieving better generalization and convergence in object representing. Our work also groups it but for scale-invariance and -variance.
Multi-Scale. Objects exist at different scales due actual sizes and imaging distances. CNN (zhang2018shufflenet) and ViT (gu2022hrvit) typically have their comfort zones as the scale distribution of objects in training data is always nonuniform. In computer vision, like detection (zou2023objectdetection) and segmentation (minaee2021imagesegmentation), the multi-scale issue is significant, leading to the wide adoption of “divide-and-rule”. The image pyramid (adelson1984imagepyramid) produces multi-scale representations by resizing; Feature pyramid network (lin2017featurepyramid) and variants mix features of different layers by channel-concat and element-wise-sum. We are the first to realize multi-scale fusion on VAE, augmenting every scale with the high-quality representation in other scales.
@@ 3 Proposed Method
We propose Multi-Scale Fusion (MSF), to explicitly address the multi-scale issue in VAE intermediate representation, and thereby to guide Object-Centric Learning (OCL) better. Our technique is applicable to both transformer-based OCL (singh2021slate; singh2022steve) classics and diffusion-based OCL (wu2023slotdiffuz) state-of-the-arts, which are our basis methods. Note that two types of methods use dVAE (im2017dvae) and VQ-VAE (van2017vqvae) respectively, but we unify them with VQ-VAE like that in GDR (zhao2024gdr).
@@ 3.1 Background: OCL with VAE Guidance
Initially, the primary encoder encodes the input into a feature map. $\bm{\phi}_{\mathrm{e}}:\bm{T}\rightarrow\bm{F}$, where $\bm{F}\in\mathbb{R}^{h\times w\times c}$, and $\bm{\phi}_{e}$ is a CNN (he2016resnet) or ViT (caron2021dino) module. Typically, $h=w=32$ and $c=256$.
Then, the Slot Attention (SA) (locatello2020slotattent) is used to aggregate the dense feature map $\bm{F}$ into sparse object-level feature vectors, i.e., slots $\bm{S}\in\mathbb{R}^{s\times c}$, along with the byproduct, i.e., segmentation masks $\bm{M}\in\mathbb{R}^{s\times h\times w}$, under query vectors $\bm{S}_{0}$. $\bm{\phi}_{\mathrm{sa}}:(\bm{S}_{0},\bm{F})\rightarrow(\bm{S},\bm{M})$, which is basically an iterative Query-Key-Value attention (bahdanau2014qkvattent) mechanism with $\bm{S}_{0}$ as the query and $\bm{F}$ as the key and value. Typically, $s$ is a predefined number, usually the maximum number of objects plus the background in an image or video sample of a dataset.
Meanwhile, pretrained VAE (im2017dvae; van2017vqvae) converts $\bm{T}$ into continuous intermediate representation. $\bm{\phi}_{\mathrm{e}}^{\mathrm{v}}:\bm{T}\rightarrow\bm{Z}$, where $\bm{Z}\in\mathbb{R}^{h\times w\times c}$ and typically $h=w=32$ and $c=256$. This continuous representation is quantized into discrete representation by matching continuous representation with codebook and selecting matched codes to form discrete representation. $f_{\mathrm{q}}=f_{\mathrm{m}}\circ f_{\mathrm{s}}:(\bm{Z},\bm{C})\rightarrow(\bm{P},\bm{I},\bm{X})$, where codebook $\bm{C}\in\mathbb{R}^{m\times c}$, matching probabilities $\bm{P}\in\mathbb{R}^{h\times w\times m}$, matched indexes $\bm{I}\in\mathbb{R}^{h\times w}$ and discrete representation $\bm{X}\in\mathbb{R}^{h\times w\times c}$. The matching $f_{\mathrm{m}}$ and selection $f_{\mathrm{s}}$ are implemented as below, respectively
[…]
@@
Lastly, with $\bm{S}$ as the condition, the primary decoder reconstructs the input guided by VAE discrete intermediate representation conditioned on $\bm{S}$, which drives every slot to aggregate as much object information as possible. $\bm{\phi}_{\mathrm{d}}:(\bm{X},\bm{S})\rightarrow\bm{X}^{\prime}$, where $\bm{X}^{\prime}\in\mathbb{R}^{h\times w\times c}$. For transformer-based OCL, a transformer decoder (vaswani2017transformer) with internal masking performs the reconstruction as classification. For diffusion-based OCL, a diffusion model, typically a conditional UNet (rombach2022ldm), carries out the reconstruction as regression of the noise added to $\bm{X}$.
@@ 3.2 Naive Multi-Scale Training
There can be objects of different sizes in images or videos, while models usually have best-at patterns scales due to non-uniform data distributions. This makes image pyramid (adelson1984imagepyramid) an effective technique, where the input is resized into different sizes and processed in parallel. Such multi-scale representations make the pattern scale distribution more uniform for training, and ensures different sized objects to fall within models’ comfort zone for testing.
Input image or video frame $\bm{T}_{1}:=\bm{T}$ is down-sampled with typical bi-linear interpolation into $\bm{T}_{n}\in\mathbb{R}^{\frac{h_{0}}{2^{n-1}}\times\frac{w_{0}}{2^{n-1}}\times c_{0}}$, where n=1Nn=1...N means different spatial scales and we use $N=3$.
@@ 3.3 Inter/Intra-Scale Quantized Fusion
Firstly, we project VAE continuous representation of a certain scale $\bm{Z}_{n}$ into two copies, $\bm{Z}_{n}^{\mathrm{si}}$ for scale-invariance and $\bm{Z}_{n}^{\mathrm{sv}}$ for scale-variance, respectively. These two will finally be combined back, thus we employ the invertible project-up-project-down design proposed in zhao2024gdr to maintain the inductive biases added in between.
@@
Meanwhile, for the scale-invariant continuous representation, we conduct inter-scale fusion over scale-invariant quantization of all scales. $f_{\mathrm{rf}}:\{(\bm{Z}_{n}^{\mathrm{si}},\bm{C}^{\mathrm{si}})\}\rightarrow\{(\dot{\bm{P}}_{n}^{\mathrm{si}},\dot{\bm{I}}_{n}^{\mathrm{si}},\dot{\bm{X}}_{n}^{\mathrm{si}})\}$, where n=1Nn=1...N. The shared codebook enforces scale-invariant representation learning.
Now that all scales $\{\bm{X}_{n}^{\mathrm{si}}\}$ consist of codes in a shared codebook $\bm{C}^{\mathrm{si}}$, matching probabilities $\{\bm{P}_{n}^{\mathrm{si}}\}$ for the same (super-)pixel region under different scales indicate how confident the model is about the representation quality. We utilize this to augment the low-quality representation in some scales with the high-quality in other scales. Take scales n1n-1, $n$ and $n+1$ as an example:
@@
    Given any super-pixel $\bm{z}_{n}^{\mathrm{si}}:=\bm{Z}_{n}^{\mathrm{si},(i,j,:)}\in\mathbb{R}^{c}$ in any scale $n$, find the corresponding region/super-pixel(s) in other scales, and mean them along all spatial dimensions as voting. For scale n1n-1, we choose $\bm{Z}_{n-1}^{\mathrm{si},(2i:2i+2,2j:2j+2,:)}\in\mathbb{R}^{2\times 2\times c}$ and mean it into a vector $\bm{z}_{n-1}^{\mathrm{si}}\in\mathbb{R}^{c}$; For scale $n+1$, we choose $\bm{z}_{n+1}^{\mathrm{si}}:=\bm{Z}_{n+1}^{\mathrm{si},(\lfloor\frac{i}{2}\rfloor,\lfloor\frac{j}{2}\rfloor,:)}\in\mathbb{R}^{c}$.
@@ 3.4 Resource Consumption Analysis
Codebook-related parameters. The basis methods use a codebook of size $(m,c)=4096\times 256$, which amounts to 1.0481.048M. We use (1) invertiable projection to project the continuous representation into two copies (and project them back), up to $c\times 2c=2^{17}$; (2) four groups of codebooks of size $(m^{0.5},c)=(64,256)$, one shared for $N=3$ scale-invariant scales, three specifically for $N=3$ scale-variant scales, up to $m^{0.5}\times c\times(1+N)=2^{16}$. Thus our codebook-related parameters are 0.1960.196M – we have $81.25\%$ fewer codebook parameters than the basis methods.
Computation and memory consumption. The exact calculation relates with the VAE encoder/decoder layers, like Conv2d, GroupNorm, Mish, etc., thus has too many details. But due to the $O(n^{2})$ complexity nature of convolution layers, we can quickly get the estimation in ratio. Denote the computation of the original VAE as unit, then our $N=3$ scales’ computations would be $1$, $\frac{1}{4}$ and $\frac{1}{16}$, since the latter two has $2\times$ and $4\times$ down sampling rates. Thus in VAE pretraining, our technique requires roughly $31.25\%$ more computation. The memory consumption increase is similar.
@@ 4.1 OCL Performance
To evaluate the quality of OCL object representations, i.e., slots, we use the accuracy of OCL’s byproduct (unsupervised) segmentation as an intuitive measurement. Metrics including ARI (Adjusted Rand Index) [^1] https://scikit-learn.org/stable/modules/generated/sklearn.metrics.adjusted_rand_score.html, ARIfg (foreground), mIoU (mean Intersection-over-Union) [^2] https://scikit-learn.org/stable/modules/generated/sklearn.metrics.jaccard_score.html and mBO (mean Best Overlap) (caron2021dino) are used to measure OCL’s byproduct segmentation accuracy. Datasets are either synthetic images, i.e., ClevrTex [^3] https://www.robots.ox.ac.uk/~vgg/data/clevrtex/; or real-world images, i.e., COCO [^4] https://cocodataset.org/#panoptic-2020 and VOC [^5] http://host.robots.ox.ac.uk/pascal/VOC/voc2012/index.html; or synthetic videos, i.e., MOVi-C/D/E [^6] https://github.com/google-research/kubric/tree/main/challenges/movi.
To evaluate how MSF boosts OCL, we use transformer-based classics, SLATE (singh2021slate) for images and STEVE (singh2022steve) for videos; use diffusion-based state-of-the-arts, SlotDiffusion (wu2023slotdiffuz) for images. The foundation-based DINOSAUR (seitzer2023dinosaur) and competitors SysBinder (singh2022sysbind) and GDR (zhao2024gdr) are compared too. All primary encoders are unified as DINO (caron2021dino) to form strong baselines.
Our MSF is a general improver to both transformer-based classics and diffusion-based state-of-the-arts, as shown in Tab. 2, 2 and 4. On complex synthetic dataset ClevrTex, MSF improves SLATE significantly, up to 12% in ARIfg. On challenging real-world dataset COCO, MSF still manages to improve SLATE by 6% in ARIfg. MSF even boosts SlotDiffusion, which is state-of-the-art, by 3% in ARI on VOC. Compared with other improvers, i.e., SysBinder@gg2 and GDR@gg2, our MSF also beats or at least ties them on SLATE, while beats them consistently on STEVE. On SlotDiffusion, our MSF beats the strong competitor GDR@gg2, and defeats DINOSUAR.
[…]